what is the phase shift for 3sin(3x + Pi/2)
A) 3
B) -3Pi/2
C) -Pi/6
D) Pi/2

Answer by stanbon(75887)

what is the phase shift for 3sin(3x + Pi/2)
In the form y=asin(bx-c) the phase shift is c/b
For you -c=pi/2 and b=3
So, c/b = -pi/2/3=-pi/6
Cheers,
Stan H.